OpenStudy (anonymous):

A box contains five light globes, two of which are faulty. Two globes are selected, one at a time without replacement. find the probability that: (a) both globes are faulty (b) neither globe is faulty (c)the first globe is faulty and the second one is not (d)the second globe is faulty and the first one is not.

OpenStudy (anonymous):

(a) (2/5)*(1/4)=2/20=1/10 (b) (3/5)*(2/4)=6/20=3/10 (c) (2/5)*(3/4)=6/20=3/10 (d) (3/5)*(2/4)=3/10

OpenStudy (anonymous):

thank you:)